13.) What is the pH of a solution if 25.0 mL of 0.250 M HClO4 is titrated with 25.0 mL of 0.475 M Ba(OH)2?

a. 2.30

b. 10.32

c. 2.73

d. 13.54

e. 7.00
Answer #1

Ba(OH)2 + 2 HClO4 ------------> Ba(ClO4)2 + 2H2O

millimoles of HClO4 = 25 x 0.25 = 6.25

millimoles of Ba(OH)2 = 25 x 0.475 = 11.875

millimoles of OH-  = 2 x 11.875 = 23.75

23.75 - 6.25 = 17.5 millimoles of OH-   left

[OH-] = 17.5 / 50 = 0.35 M

pOH = - log [OH-]

pOH = - log [0.35]

pOH = 0.45

pH = 14 - 0.45

pH = 13.55

answer = option d = 13.54